OpenStudy (anonymous):

What constant acceleration is required to increase the speed of a car from 23 mi/h to 55 mi/h in 3 s? (Round your answer to two decimal places.)

OpenStudy (anonymous):

55-23/3=10.66 m/s^2

OpenStudy (anonymous):

im sorry i need it in ft/s^2

OpenStudy (anonymous):

divide the change in velocities with the time taken

OpenStudy (anonymous):

do you how to change ft to meters?

OpenStudy (anonymous):

v=u+at v-u=at \[a=\frac{ v-u }{ t }\]

OpenStudy (anonymous):

i need feet. i have to convert feet to meters.?

OpenStudy (anonymous):

55m/h=\[55m/hr=\frac{ 55*1760*3 }{ 1*60*60 }=? ft/\sec\]

OpenStudy (anonymous):

80.6666 repeating 6's

OpenStudy (anonymous):

1 mile=1760 yards 1 yard=3ft 1hr=60 min 1min=60 sec.

OpenStudy (anonymous):

Thank you!